# Anomaly Prediction: A Novel Approach with Explicit Delay and Horizon

### v2

@@ III-B3 Lead Time
$\text{LeadTime}(\hat{Y}_{t},Y_{t})=1-\frac{1}{T}\left|i_{\hat{y}}-i_{y}\right|$
where $i_{y}=\min\{i\mid y_{t+i}=1\}$ is the index of the first observed ground truth anomaly, $i_{\hat{y}}=\min\{i\mid\hat{y}_{t+i}>s\}$ is the index of the first predicted anomaly probabilities that are greater than user-defined threshold $s$.
This metric ranges from 0 to 1, where 1 indicates that the first predicted anomaly exactly matches the timing of the first ground truth anomaly, and values closer to 0 indicate a larger discrepancy between the predicted and actual timing.
@@ III-B4 Dice Score
$\text{DiceScore}(\hat{Y}_{t},Y_{t})=\frac{2|\hat{Y}_{t}\cap Y_{t}|}{|\hat{Y}_{t}|+|Y_{t}|}$
This metric ranges from 0 to 1, where 1 indicates perfect coverness between the predicted anomalies $\hat{y}$ and the ground truth $y$, and 0 indicates no overlap.
@@ III-C Models
In this section, we describe the models employed in our experiments for anomaly detection in time series data. Our selection includes some of the most recent and fondamental techniques: FCN, PatchTST, Kernel U-Net. These models were chosen for their simplicity and demonstrated performance in the field of time series forecasting.
Fully Connected Network Fully Connected Network (FCN) serves as a baseline model for anomaly prediction in time series data. This model applies multiple linear layers with ReLU activation functions to automatically extract temporal patterns. Due to its simplicity and effectiveness, we use it as a baseline to compare the performance of more complex models in time series anomaly prediction.
PatchTST (Patch Time Series Transformer) [51] is a novel model designed to capture long-term dependencies in time series data through the use of patches, analogous to those used in computer vision transformers. The core idea of PatchTST is to segment the time series into patches and process them using transformer encoders. This approach enables the model to efficiently learn and represent temporal patterns over extended periods. PatchTST leverages self-attention mechanisms to weigh the importance of different time steps, making it particularly effective in identifying anomalies that have complex temporal dependencies.
Kernel U-Net [38] is an extension of the traditional U-Net architecture, specifically designed to address the challenges in time series forecasting. Traditional Transformer-based U-Net models, while successful in medical image segmentation, often struggle with expressiveness and computational efficiency in time series applications. Kernel U-Net overcomes these limitations through a symmetric, hierarchical U-shaped neural network structure. Its encoder compresses input time series into latent vectors, which are then expanded into output series by the symmetric decoder. By separating the process of partitioning input time series into patches from kernel manipulation, Kernel U-Net allows for customized kernel execution, enhancing flexibility. This approach not only adapts to specific datasets but also improves computational efficiency, reducing the complexity of the Transformer layer to linear.
@@ IV Dataset
Real-world datasets for time series anomaly prediction can be scarce, incomplete, or lack sufficient anomalies to train and evaluate models effectively. This limitation necessitates the creation of synthetic datasets that can simulate various scenarios of anomalies. A synthetic dataset allows researchers to control the parameters and conditions under which anomalies occur, providing a robust platform for benchmarking and validating anomaly prediction models. In this section, we describe a synthetic dataset, a private dataset designed for anomaly prediction, and 3 datasets from anomaly detection. The descriptions of datasets are in Table II.
@@ IV-A1 Synthetic Time Series Anomaly
  • Time Series Pattern: A series of points generated by sine or cosine function. We use single fixed sine function (Fixed), mixed sine function of different frequencies (Mixed), multivariate series composed of sine or cosine functions (Mixed).
  • Anomaly Pattern: A pattern that the model can detect. It is usually not observable by human experts but contains decisive patterns indicative of an anomaly in the future. We use a fixed pattern composed of a segment of gaussian random points and a increasing line (Fixed).
@@
  • Observation: A sequence of annotation of anomaly after the brewing time. It is typically visually evident to human experts, prompting them to raise an alert. The anomaly event is labeled as a positive example during the observation. We use a fixed number (Fixed) or a random number following gaussian distribution (Gaussian) for the size of anomaly labels.
Table III resumes the details of generated synthetic dataset for anomaly prediction.
@@ IV-A2 Signal-to-Noise Ratio (SNR)
Synthetic_1 0.001 0.001 0.001 0.956 0.643 0.958 0.995 0.994 0.996 0.936 0.852 0.953 0.391 0.272 0.369
Synthetic_2 0.002 0.002 0.002 0.783 0.532 0.873 0.995 0.995 0.995 0.879 0.847 0.884 0.131 0.093 0.143
Synthetic_3 0.005 0.006 0.005 0.956 0.898 0.963 0.991 0.990 0.991 0.925 0.904 0.935 0.700 0.669 0.710
Synthetic_4 0.008 0.010 0.007 0.901 0.772 0.906 0.987 0.987 0.988 0.885 0.853 0.904 0.588 0.566 0.588
Synthetic_5 0.011 0.011 0.011 0.897 0.788 0.891 0.986 0.986 0.986 0.808 0.795 0.827 0.508 0.498 0.517
Synthetic_6 0.001 0.001 0.001 0.935 0.258 0.960 0.996 0.995 0.996 0.943 0.840 0.950 0.406 0.115 0.459
Synthetic_7 0.002 0.002 0.001 0.785 0.024 0.852 0.995 0.996 0.995 0.833 0.850 0.868 0.148 0.000 0.173
Synthetic_8 0.012 0.010 0.010 0.912 0.469 0.947 0.988 0.985 0.990 0.911 0.785 0.927 0.676 0.375 0.692
Synthetic_9 0.019 0.019 0.017 0.826 0.305 0.838 0.985 0.982 0.985 0.805 0.720 0.807 0.507 0.392 0.516
Synthetic_10 0.019 0.021 0.020 0.465 0.149 0.786 0.982 0.986 0.985 0.815 0.514 0.823 0.392 0.062 0.482
SBDA 0.087 0.109 0.090 0.398 0.396 0.352 0.997 0.995 0.997 0.741 0.754 0.739 0.158 0.115 0.182
@@ IV-B1 SMD
IV-B2 SBDA
The server breakdown anomaly (SBDA) dataset is a private anonymized dataset. It is composed of 16 metrics, such as CPU, memory usage, and other system-related measurements, and 36 label columns annotated by experts. It contains 10 data instances collected over six months, capturing normal operations and server failures. This dataset is instrumental for evaluating and developing anomaly detection algorithms to identify server breakdowns. Compared with the SMD dataset, the SBDA dataset contains not only physical metrics but also additional expert knowledge through annotations.
@@ V Experiments and Results
The experimental results demonstrated the high accuracy of the models in detecting the presence of anomalies across different synthetic datasets (Table IV). FCN and K-U-Net correctly identified anomalies in over 90% of the test cases in a univariate setting, as indicated by the existence of anomaly metric. However, PatchTST quickly overfitted to particular cases, achieving only a 75% F1-score. We measured the density sum, lead time, and Dice score in cases where the models successfully predicted the existence of an anomaly (true positives). The density sum showed that the models’ predictions closely matched the total anomalies expected in the future horizon, with all models reaching 0.95. The lead time metric indicated that the models’ predictions were closely aligned with the actual occurrences of anomalies, with an average lead time score of 0.85, suggesting effective prediction with minimal delay. The Dice score averaged below 0.5 due to the challenges posed by temporal dynamics, reflecting that the overlap between predicted and actual anomalies varied depending on the distribution of anomaly labels and brewing time. Results for the existence of anomaly on real-world datasets such as SBDA, SMD, SMAP, and MSL were generally lower than 0.5.
Overall, these results provide proof of concept for our anomaly prediction approach. The use of synthetic datasets allowed us to rigorously test the models under controlled conditions, ensuring their reliability before applying them to real-world applications. We share examples of anomaly prediction in Figure 4, 5. The results on real-world datasets initially validate the concept but also reveal existing difficulties in achieving high accuracy.
@@ VI Conclusion
This work advances the field of anomaly detection by proposing the novel approach of anomaly prediction, laying a solid foundation for future research and practical applications. Future work may focus on enhancing the model’s adaptability and precision, addressing the identified challenges such as accuracy and temporal dynamics, and developing more reliable and effective anomaly prediction tools for diverse real-world applications.

### v3

@@ III-B3 Lead Time
$\text{LeadTime}(\hat{Y}_{t},Y_{t})=1-\frac{1}{T}\left|i_{\hat{y}}-i_{y}\right|$ (8)
where $i_{y}=\min\{i\mid y_{t+i}=1\}$ is the index of the first observed ground truth anomaly, $i_{\hat{y}}=\min\{i\mid\hat{y}_{t+i}\geq s\}$ is the index of the first predicted anomaly probabilities that are greater than user-defined threshold $s$.
This metric ranges from 00 to $1$, where $1$ indicates that the first predicted anomaly exactly matches the timing of the first ground truth anomaly, and values closer to 0 indicate a larger discrepancy between the predicted and actual timing.
@@ III-B4 Dice Score
$\text{DiceScore}(\hat{Y}_{t},Y_{t})=\frac{2|\hat{Y}_{t}^{\prime}\cap Y_{t}|}{|\hat{Y}_{t}^{\prime}|+|Y_{t}|}$ (9)
where $\hat{Y}_{t}^{\prime}=\mathbbm{1}(\hat{Y}_{t}\geq s)$ is an indicator function that converts the probabilities to booleans. This metric ranges from 00 to $1$, where $1$ indicates perfect coverness between the predicted anomalies $\hat{y}$ and the ground truth $y$, and 00 indicates no overlap.
@@ III-C Models
In this section, we describe the models employed in our experiments for anomaly detection in time series data. Our selection includes some of the most recent and fundamental techniques: Fully Connected Network, PatchTST, and Kernel U-Net. These models were chosen for their simplicity and demonstrated performance in the field of time series forecasting.
Fully Connected Network (FCN) serves as a baseline model for anomaly prediction in time series data. This model applies multiple linear layers with ReLU activation functions to extract temporal patterns. Due to its simplicity and effectiveness, we use it as a baseline to compare the performance of more complex models in time series anomaly prediction.
PatchTST (Patch Time Series Transformer) [47] is a novel model designed to capture long-term dependencies in time series data through the use of patches, analogous to those used in computer vision transformers. The core idea of PatchTST is to segment the time series into patches and process them using transformer encoders. This approach relieves the overfitting of the transformer and reduces the complexity to $O(\frac{L}{S}log(\frac{L}{S}))$, where $L$ is the length and $S$ is patch size.
Kernel U-Net [34] is an extension of the traditional U-Net architecture, specifically designed to address the challenges in time series forecasting. This model separates the process of partitioning input time series into patches from kernel manipulation, allowing for customized kernel execution. Kernel U-Net keeps a symmetric, hierarchical U-shaped neural network structure, which ensures its linear computation complexity $O(L)$, and makes it more suitable for diverse and large-scale time series applications.
@@ IV Dataset
Real-world datasets for time series anomaly prediction can be scarce, incomplete, or lack sufficient anomalies to train and evaluate models effectively. This limitation necessitates the creation of synthetic datasets that can simulate various scenarios of anomalies. Synthetic datasets allow researchers to control the parameters and conditions under which anomalies occur, providing a robust platform for benchmarking and validating anomaly prediction models. In this section, we describe a synthetic dataset, a private dataset designed for anomaly prediction, and 3 datasets from anomaly detection. The descriptions of datasets are in Table II.
@@ IV-A1 Synthetic Time Series Anomaly
  • Time Series Pattern: A series of points generated by sine or cosine function. We use a single fixed sine function (Fixed), a mixed sine function of different frequencies (Mixed), multivariate series composed of sine or cosine functions (Mixed).
  • Anomaly Pattern: A pattern that the model can detect. It is usually not observable by human experts but contains decisive patterns indicative of an anomaly in the future. We use a fixed pattern composed of a segment of Gaussian random points and an increasing line (Fixed).
@@
  • Observation: A sequence of annotation of anomaly after the brewing time. It is typically visually evident to human experts, prompting them to raise an alert. The anomaly event is labeled as a positive example during the observation. We use a fixed number (Fixed) or a random number (Gaussian) for the size of anomaly labels.
Table III resumes the details of the generated synthetic dataset for anomaly prediction.
@@ IV-A2 Signal-to-Noise Ratio (SNR)
Synthetic_1 0.003 0.003 0.004 0.956 0.643 0.958 0.995 0.994 0.996 0.936 0.852 0.953 0.391 0.272 0.369
Synthetic_2 0.004 0.005 0.004 0.783 0.532 0.873 0.995 0.995 0.995 0.879 0.847 0.884 0.131 0.093 0.143
Synthetic_3 0.013 0.014 0.012 0.956 0.898 0.963 0.991 0.990 0.991 0.925 0.904 0.935 0.700 0.669 0.710
Synthetic_4 0.019 0.024 0.018 0.901 0.772 0.906 0.987 0.987 0.988 0.885 0.853 0.904 0.588 0.566 0.588
Synthetic_5 0.027 0.026 0.028 0.897 0.788 0.891 0.986 0.986 0.986 0.808 0.795 0.827 0.508 0.498 0.517
Synthetic_6 0.002 0.003 0.002 0.935 0.258 0.960 0.996 0.995 0.996 0.943 0.840 0.950 0.406 0.115 0.459
Synthetic_7 0.004 0.004 0.003 0.785 0.024 0.852 0.995 0.996 0.995 0.833 0.850 0.868 0.148 0.000 0.173
Synthetic_8 0.028 0.023 0.024 0.912 0.469 0.947 0.988 0.985 0.990 0.911 0.785 0.927 0.676 0.375 0.692
Synthetic_9 0.046 0.046 0.042 0.826 0.305 0.838 0.985 0.982 0.985 0.805 0.720 0.807 0.507 0.392 0.516
Synthetic_10 0.047 0.051 0.048 0.465 0.149 0.786 0.982 0.986 0.985 0.815 0.514 0.823 0.392 0.062 0.482
SBAD 0.036 0.045 0.037 0.398 0.396 0.352 0.997 0.995 0.997 0.741 0.754 0.739 0.158 0.115 0.182
@@ IV-B1 SMD
IV-B2 SBAD
The Server Breakdown Anomaly Detection (SBAD) dataset is a private anonymized dataset. It is composed of 16 metrics, such as CPU, memory usage, and other system-related measurements, and 36 label columns annotated by experts. It contains 10 data instances collected over six months, capturing normal operations and server failures. This dataset is instrumental for evaluating and developing anomaly detection algorithms to identify server breakdowns. Compared with the SMD dataset, the SBAD dataset contains not only physical metrics but also additional expert knowledge through annotations.
@@ V Experiments and Results
The experimental results demonstrated the high accuracy of the models in detecting the presence of anomalies across different synthetic datasets (Table IV). FCN and K-U-Net correctly identified anomalies in over 90% of the test cases in a univariate setting, as indicated by the existence of anomaly metric. However, PatchTST quickly overfitted to particular cases, achieving only a 75% F1-score. We measured the Density, Lead Time, and Dice Score in cases where the models successfully predicted the existence of an anomaly (True Positives). The density sum showed that the models’ predictions closely matched the total anomalies expected in the future horizon, with all models reaching 0.95. The lead time metric indicated that the models’ predictions were closely aligned with the actual occurrences of anomalies, with an average lead time score of 0.85, suggesting effective prediction with minimal delay. The Dice score averaged below 0.5 due to the challenges posed by temporal dynamics, reflecting that the overlap between predicted and actual anomalies varied depending on the distribution of anomaly labels and brewing time. Results for the existence of anomaly on real-world datasets such as SBAD, SMD, SMAP, and MSL were generally lower than 0.5.
@@
Overall, these results provide proof of concept for our anomaly prediction approach. The use of synthetic datasets allowed us to rigorously test the models under controlled conditions, ensuring their reliability before applying them to real-world applications. The results on real-world datasets initially validate the concept but also reveal existing challenges in achieving high accuracy. We share examples of anomaly prediction on synthetic datasets in Figures 4, 5.
@@ VI Conclusion
This work advances the field of anomaly detection by proposing a novel approach to anomaly prediction, laying a solid foundation for future research and practical applications. Future work may focus on enhancing the model’s adaptability and precision, addressing the identified challenges such as accuracy and temporal dynamics, and developing more reliable anomaly prediction tools for diverse real-world applications.